OpenStudy (anonymous):

which correctly describes how the graph of inequality 4x -2y > 20 is shaded? A. Above the solid line B. Below the solid line C. Above the dashed line D.Below the dashed line

OpenStudy (anonymous):

@pooja195 @Hero

OpenStudy (anonymous):

Someone please help ;-;

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

Please no direct answers, Just want to know the steps so I can do the rest on my own

OpenStudy (e.mccormick):

Well, one way to see this more easily is to get it into something like y=mx+b form so you will know a little about the line. Then test some.

Elsa213 (elsa213):

ye.... wat he said. c:

OpenStudy (anonymous):

so you have to play around with it correct?

OpenStudy (e.mccormick):

I'll do the simple example, y<x+1. You could be given that as: y-x<1 And yes, it is manipulating it to get it into the form where it is clear what the intercept point is.

OpenStudy (e.mccormick):

On y<x+1 The intercept is 1. So when x is 0, y is 1... BUT, it is an inequality. That means y is not really 1. It is less than 1 in this case. So that tells me that wherever the line is, the shading for the y values must be below the line. The line would also be dashed because it is is < and not \(\le\).

OpenStudy (anonymous):

so is there any chance we can start you know process of elimination or we cant?

OpenStudy (e.mccormick):

Once you have it in that form, it is pretty easy to tell if it will be above or below. Just need to be careful if you do any division by a negative because that changes < into > or > into <.

OpenStudy (e.mccormick):

\(\ge\) and \(\le\) are solid lines. > and < are dashed lines. That alone elimiantes 2 answers.

OpenStudy (anonymous):

so it's solid?

OpenStudy (anonymous):

wait i mean dashed

OpenStudy (e.mccormick):

Yah. 4x -2y > 20 uses >. Even if it flips direction it would be <. So it must be dashed. When you get it into y=mx+b form you know the other part: y>mx+b means y is greater (above) the line y<mx+b means y is lesser (below) the line

OpenStudy (anonymous):

so it has to be C or D

OpenStudy (e.mccormick):

Yes.

OpenStudy (anonymous):

Think your trying to say it's D correct?

Elsa213 (elsa213):

he gone.....

OpenStudy (anonymous):

@e.mccormick @ganeshie8 @EclipsedStar

Elsa213 (elsa213):

chico..... ur correct c:

OpenStudy (anonymous):

really? :D

Elsa213 (elsa213):

eh ye.... ig

OpenStudy (e.mccormick):

Yes, I was gone... the joys of moderaton. 4x -2y > 20 -2y > -4x + 20 y < 2x - 10 So yes, below the line.

OpenStudy (anonymous):

lol xD thank you so much

OpenStudy (anonymous):

With your helpful comments now I understand